# Semi-Supervised Deep Sobolev Regression: Estimation, Variable Selection and Beyond

### v1

@@ D.1 Deep Sobolev Regressor
$\lambda(\nabla\widehat{f}_{{\mathcal{D}}}^{\lambda},\nabla(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f))_{L^{2}(\nu_{X})}$ $=\lambda(\nabla(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0})+\nabla f_{0},\nabla(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0})-\nabla(f-f_{0}))_{L^{2}(\nu_{X})}$
$=\lambda\|\nabla(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0})\|_{L^{2}(\nu_{X})}^{2}-\lambda(\nabla(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0}),\nabla(f-f_{0}))_{L^{2}(\nu_{X})}$
$\quad+\lambda(\nabla f_{0},\nabla(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f))_{L^{2}(\nu_{X})}.$ (43)
@@
$=\lambda(\Delta f_{0},\widehat{f}_{{\mathcal{D}}}^{\lambda}-f)_{L^{2}(\nu_{X})}+\lambda(\nabla f_{0}\cdot\nabla(\log q),\widehat{f}_{{\mathcal{D}}}^{\lambda}-f)_{L^{2}(\nu_{X})}$
$\leq\lambda\kappa^{1/2}\Big{\{}\|\Delta f_{0}\|_{L^{2}(\nu_{X})}+\|\nabla f_{0}\cdot\nabla(\log q)\|_{L^{2}(\nu_{X})}\Big{\}}\Big{\{}R^{1/2}(\widehat{f}_{{\mathcal{D}}}^{\lambda})+\|f-f_{0}\|_{L^{2}(\mu_{X})}\Big{\}}$
$\leq 9\lambda^{2}\kappa\Big{\{}\|\Delta f_{0}\|_{L^{2}(\nu_{X})}^{2}+\|\nabla f_{0}\cdot\nabla(\log q)\|_{L^{2}(\nu_{X})}^{2}\Big{\}}+\frac{1}{16}R(\widehat{f}_{{\mathcal{D}}}^{\lambda})+\frac{1}{2}\|f-f_{0}\|_{L^{2}(\mu_{X})}^{2},$ (44)
@@
$\lambda(\nabla(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0}),\nabla(f-f_{0}))_{L^{2}(\nu_{X})}\leq\frac{\lambda}{2}\|\nabla(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0})\|_{L^{2}(\nu_{X})}^{2}+\frac{\lambda}{2}\|\nabla(f-f_{0})\|_{L^{2}(\nu_{X})}^{2}.$ (45)
@@
$\frac{\lambda}{2}\|\nabla(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0})\|_{L^{2}(\nu_{X})}^{2}$ $\leq\lambda(\nabla\widehat{f}_{{\mathcal{D}}}^{\lambda},\nabla(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f))_{L^{2}(\nu_{X})}+\frac{1}{16}R(\widehat{f}_{{\mathcal{D}}}^{\lambda})$ (46)
@@
$(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0},\widehat{f}_{{\mathcal{D}}}^{\lambda}-f)_{L^{2}({\mathcal{D}})}$ $=\widehat{R}_{{\mathcal{D}}}(\widehat{f}_{{\mathcal{D}}}^{\lambda})-(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0},f-f_{0})_{L^{2}({\mathcal{D}})}$
@@
$\leq\frac{1}{16}{\mathbb{E}}_{{\mathcal{D}}}\Big{[}R(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}+{\mathbb{E}}_{{\mathcal{D}}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}\widehat{f}_{{\mathcal{D}}}^{\lambda}(X_{i})\Big{]}+\Big{\{}\|f-f_{0}\|_{L^{2}(\mu_{X})}^{2}+\frac{\lambda}{2}\|\nabla(f-f_{0})\|_{L^{2}(\nu_{X})}^{2}\Big{\}}$
@@
${\mathbb{E}}_{{\mathcal{D}}}\Big{[}\widehat{R}_{{\mathcal{D}}}(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}\leq\frac{1}{4}{\mathbb{E}}_{{\mathcal{D}}}\Big{[}R(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}+c\Big{\{}\beta\lambda^{2}+\varepsilon_{\operatorname{app}}({\mathcal{F}},\lambda)+\varepsilon_{\operatorname{gen}}({\mathcal{F}},n)\Big{\}},$
@@
${\mathbb{E}}_{{\mathcal{D}}}\Big{[}R(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}\leq 2{\mathbb{E}}_{{\mathcal{D}}}\Big{[}\widehat{R}_{{\mathcal{D}}}(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}+c^{\prime}\varepsilon_{\operatorname{gen}}({\mathcal{F}},n)$ (48)
$\leq\frac{1}{2}{\mathbb{E}}_{{\mathcal{D}}}\Big{[}R(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}+2c\Big{\{}\beta\lambda^{2}+\varepsilon_{\operatorname{app}}({\mathcal{F}},\lambda)\Big{\}}+(2c+c^{\prime})\varepsilon_{\operatorname{gen}}({\mathcal{F}},n).$
@@
$\|f-\phi\|_{L^{2}(\mu_{X})}\leq CN^{-s}\quad\text{and}\quad\|\nabla(f-\phi)\|_{L^{2}(\nu_{X})}\leq CN^{-(s-1)}.$
By using Lemma A.6, we find $\operatorname{VCdim}({\mathcal{F}})\leq cN^{d}\log N$. Substituting these estimates into Lemma 5.2 yields
@@
$\lesssim\beta\lambda^{2}+CN^{-2s}+C\lambda N^{-2(s-1)}+(B_{0}^{2}+\sigma^{2})\frac{N^{d}(\log N)^{2}}{n\delta}+(B_{0}+\sigma)\delta.$
Setting $N=(1+\sigma)^{-\frac{3}{d+4s}}n^{\frac{1}{d+4s}}$, $\delta=\sigma^{\frac{1}{2}}N^{\frac{d}{2}}n^{-\frac{1}{2}}$ and letting the regularization parameter be $\lambda=(1+\sigma)^{\frac{3s}{d+4s}}n^{-\frac{s}{d+4s}}$ deduce the desired result. ∎
D.2 Semi-Supervised Deep Sobolev Regressor
Proof of Lemma 5.5.
@@ D.2 Semi-Supervised Deep Sobolev Regressor
$(u,v)_{L^{2}({\mathcal{S}})}=\frac{1}{m}\sum_{i=1}^{m}u(Z_{i})v(Z_{i}),\quad\|u\|_{L^{2}({\mathcal{S}})}^{2}=\frac{1}{m}\sum_{i=1}^{m}u^{2}(Z_{i}),\quad u,v\in L^{\infty}(\nu_{X}),$
The proof is divided into four parts which are denoted by (I) to (IV):
@@
    By a same argument as (I) in the proof of Lemma 5.2, we deduces
    ${\mathbb{E}}_{{\mathcal{S}}}\Big{[}\|\nabla(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0})\|_{L^{2}(\nu_{X})}^{2}-2\|\nabla(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0})\|_{L^{2}({\mathcal{S}})}^{2}\Big{]}\leq c\varepsilon_{\operatorname{gen}}^{\operatorname{reg}}({\mathcal{F}},m),$ (49)
    where $c$ is an absolute positive constant. Here the generalization error $\varepsilon_{\operatorname{gen}}^{\operatorname{reg}}({\mathcal{F}},m)$ associated to the regularization term are defined as
    $\varepsilon_{\operatorname{gen}}^{\operatorname{reg}}({\mathcal{F}},m)=\inf_{\delta>0}\Big{\{}B_{1}^{2}\frac{\operatorname{VCdim}(\nabla{\mathcal{F}})}{m\delta}+B_{1}\delta\Big{\}}.$
@@
    $-\lambda{\mathbb{E}}_{{\mathcal{S}}}\Big{[}(\nabla f_{0},\nabla(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f))_{L^{2}({\mathcal{S}})}\Big{]}\leq\frac{1}{16}{\mathbb{E}}_{{\mathcal{S}}}\Big{[}R(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda})\Big{]}+c\Big{\{}\tilde{\beta}\lambda^{2}+\varepsilon_{\operatorname{gen}}^{\operatorname{reg}}({\mathcal{F}},m)\Big{\}},$ (50)
@@
    $\lambda{\mathbb{E}}_{{\mathcal{D}},{\mathcal{S}}}\Big{[}\|\nabla(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0})\|_{L^{2}({\mathcal{S}})}^{2}\Big{]}+{\mathbb{E}}_{{\mathcal{D}},{\mathcal{S}}}\Big{[}\widehat{R}_{{\mathcal{D}}}(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda})\Big{]}$ (51)
@@
    $\quad+c\Big{\{}\tilde{\beta}\lambda^{2}+\varepsilon_{\operatorname{app}}({\mathcal{F}},\lambda)+\varepsilon_{\operatorname{gen}}^{\operatorname{reg}}({\mathcal{F}},m)\Big{\}},$
@@
Step (I). By a same argument as Step (I) in the proof of Lemma 5.2, we deduce the following inequality
${\mathbb{E}}_{{\mathcal{S}}}\Big{[}\|D_{k}(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0})\|_{L^{2}(\nu_{X})}^{2}-2\|\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0}\|_{L^{2}({\mathcal{S}})}^{2}\Big{]}\leq c\inf_{\delta>0}\Big{\{}B_{1,k}^{2}\frac{\operatorname{VCdim}(D_{k}{\mathcal{F}})}{m\delta}+B_{1,k}\delta\Big{\}},$
for each $1\leq k\leq d$, where $c>0$ is an absolute constant. Summing over these equalities obtains (49) immediately.
@@
$={\mathbb{E}}_{{\mathcal{S}}}\Big{[}\sup_{f\in\operatorname{conv}({\mathcal{F}})}{\mathbb{E}}_{{\mathcal{S}}^{\prime}}\Big{[}\frac{1}{m}\sum_{i=1}^{m}D_{k}f_{0}(Z_{i}^{\prime})D_{k}f(Z_{i}^{\prime})\Big{]}-\frac{1}{m}\sum_{i=1}^{m}D_{k}f_{0}(Z_{i})D_{k}f(Z_{i})\Big{]}$
$\leq{\mathbb{E}}_{{\mathcal{S}}}{\mathbb{E}}_{{\mathcal{S}}^{\prime}}\Big{[}\sup_{f\in\operatorname{conv}({\mathcal{F}})}\frac{1}{m}\sum_{i=1}^{m}D_{k}f_{0}(Z_{i}^{\prime})D_{k}f(Z_{i}^{\prime})-\frac{1}{m}\sum_{i=1}^{m}D_{k}f_{0}(Z_{i})D_{k}f(Z_{i})\Big{]}$
$={\mathbb{E}}_{{\mathcal{S}}}{\mathbb{E}}_{{\mathcal{S}}^{\prime}}{\mathbb{E}}_{\varepsilon}\Big{[}\sup_{f\in\operatorname{conv}({\mathcal{F}})}\frac{1}{m}\sum_{i=1}^{m}\varepsilon_{i}\Big{(}D_{k}f_{0}(Z_{i}^{\prime})D_{k}f(Z_{i}^{\prime})-D_{k}f_{0}(Z_{i})D_{k}f(Z_{i})\Big{)}\Big{]}$
$={\mathbb{E}}_{{\mathcal{S}}}{\mathbb{E}}_{\varepsilon}\Big{[}\sup_{f\in\operatorname{conv}({\mathcal{F}})}\frac{1}{m}\sum_{i=1}^{m}\varepsilon_{i}D_{k}f_{0}(Z_{i})D_{k}f(Z_{i})\Big{]},$ (52)
where the second inequality follows from the Jensen’s inequality. According to Ledoux-Talagrand contraction inequality (Mohri, Rostamizadeh and Talwalkar, 2018, Lemma 5.7), we have
${\mathbb{E}}_{\varepsilon}\Big{[}\sup_{f\in\operatorname{conv}({\mathcal{F}})}\frac{1}{m}\sum_{i=1}^{m}\varepsilon_{i}D_{k}f_{0}(Z_{i})D_{k}f(Z_{i})\Big{]}\leq B_{1,k}{\mathbb{E}}_{\varepsilon}\Big{[}\sup_{f\in\operatorname{conv}({\mathcal{F}})}\frac{1}{m}\sum_{i=1}^{m}\varepsilon_{i}D_{k}f(Z_{i})\Big{]}.$ (53)
Let $\delta>0$ and let $\operatorname{conv}(D_{k}{\mathcal{F}})_{\delta}$ be an $L^{2}({\mathcal{S}})$ $\delta$-cover of $\operatorname{conv}(D_{k}{\mathcal{F}})$. Suppose $|\operatorname{conv}(D_{k}{\mathcal{F}})_{\delta}|=N(\delta,\operatorname{conv}(D_{k}{\mathcal{F}}),L^{2}({\mathcal{S}}))$. Then it follows from Cauchy-Schwarz inequality that for each $D_{k}f\in\operatorname{conv}(D_{k}{\mathcal{F}})$, there exists $(D_{k}f)_{\delta}\in\operatorname{conv}(D_{k}{\mathcal{F}})_{\delta}$ such that
@@
$\leq\Big{(}\frac{1}{m}\sum_{i=1}^{m}\varepsilon_{i}^{2}\Big{)}^{1/2}\Big{(}\frac{1}{m}\sum_{i=1}^{m}(D_{k}f(Z_{i})-(D_{k}f)_{\delta}(Z_{i}))^{2}\Big{)}^{1/2}\leq\delta,$
@@
${\mathbb{E}}_{{\mathcal{S}}}{\mathbb{E}}_{\varepsilon}\Big{[}\sup_{f\in\operatorname{conv}({\mathcal{F}})}\frac{1}{m}\sum_{i=1}^{m}\varepsilon_{i}D_{k}f(Z_{i})\Big{]}$ $\leq{\mathbb{E}}_{{\mathcal{S}}}{\mathbb{E}}_{\varepsilon}\Big{[}\sup_{f\in\operatorname{conv}({\mathcal{F}})_{\delta}}\frac{1}{m}\sum_{i=1}^{m}\varepsilon_{i}D_{k}f(Z_{i})\Big{]}+\delta$
$\leq B_{1,k}\Big{(}\frac{2\log N(\delta,\operatorname{conv}(D_{k}{\mathcal{F}}),L^{2}({\mathcal{S}}))}{m}\Big{)}^{1/2}+\delta,$ (54)
where the last inequality holds from Massart’s lemma (Mohri, Rostamizadeh and Talwalkar, 2018, Theroem 3.7). Combining (52), (53) and (54) deduces

### v2

@@ D-A Deep Sobolev Regressor
$\lambda(\nabla\widehat{f}_{{\mathcal{D}}}^{\lambda},\nabla(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f))_{L^{2}(\nu_{X})}$
$=\lambda(\nabla(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0})+\nabla f_{0},\nabla(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0})-\nabla(f-f_{0}))_{L^{2}(\nu_{X})}$
$=\lambda\|\nabla(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0})\|_{L^{2}(\nu_{X})}^{2}+\lambda(\nabla f_{0},\nabla(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f))_{L^{2}(\nu_{X})}$
$\quad-\lambda(\nabla(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0}),\nabla(f-f_{0}))_{L^{2}(\nu_{X})}.$ (49)
@@
$=\lambda(\Delta f_{0},\widehat{f}_{{\mathcal{D}}}^{\lambda}-f)_{L^{2}(\nu_{X})}$
$\quad+\lambda(\nabla f_{0}\cdot\nabla(\log q),\widehat{f}_{{\mathcal{D}}}^{\lambda}-f)_{L^{2}(\nu_{X})}$
$\leq\lambda\kappa^{1/2}\Big{\{}\|\Delta f_{0}\|_{L^{2}(\nu_{X})}+\|\nabla f_{0}\cdot\nabla(\log q)\|_{L^{2}(\nu_{X})}\Big{\}}$
×$\quad\times\Big{\{}R^{1/2}(\widehat{f}_{{\mathcal{D}}}^{\lambda})+\|f-f_{0}\|_{L^{2}(\mu_{X})}\Big{\}}$
$\leq 9\lambda^{2}\kappa\Big{\{}\|\Delta f_{0}\|_{L^{2}(\nu_{X})}^{2}+\|\nabla f_{0}\cdot\nabla(\log q)\|_{L^{2}(\nu_{X})}^{2}\Big{\}}$
$\quad+\frac{1}{16}R(\widehat{f}_{{\mathcal{D}}}^{\lambda})+\frac{1}{2}\|f-f_{0}\|_{L^{2}(\mu_{X})}^{2},$ (50)
@@
$\lambda(\nabla(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0}),\nabla(f-f_{0}))_{L^{2}(\nu_{X})}$ (51)
$\leq\frac{\lambda}{2}\|\nabla(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0})\|_{L^{2}(\nu_{X})}^{2}+\frac{\lambda}{2}\|\nabla(f-f_{0})\|_{L^{2}(\nu_{X})}^{2}.$
@@
$\frac{\lambda}{2}\|\nabla(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0})\|_{L^{2}(\nu_{X})}^{2}$ (52)
$\leq\lambda(\nabla\widehat{f}_{{\mathcal{D}}}^{\lambda},\nabla(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f))_{L^{2}(\nu_{X})}+\frac{1}{16}R(\widehat{f}_{{\mathcal{D}}}^{\lambda})$
@@
$(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0},\widehat{f}_{{\mathcal{D}}}^{\lambda}-f)_{L^{2}({\mathcal{D}})}$
$=\widehat{R}_{{\mathcal{D}}}(\widehat{f}_{{\mathcal{D}}}^{\lambda})-(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0},f-f_{0})_{L^{2}({\mathcal{D}})}$
@@
$\leq\frac{1}{16}{\mathbb{E}}_{{\mathcal{D}}}\Big{[}R(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}+{\mathbb{E}}_{{\mathcal{D}}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}\widehat{f}_{{\mathcal{D}}}^{\lambda}(X_{i})\Big{]}$
+$\quad+\Big{\{}\|f-f_{0}\|_{L^{2}(\mu_{X})}^{2}+\frac{\lambda}{2}\|\nabla(f-f_{0})\|_{L^{2}(\nu_{X})}^{2}\Big{\}}$
@@
${\mathbb{E}}_{{\mathcal{D}}}\Big{[}\widehat{R}_{{\mathcal{D}}}(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}$
$\leq\frac{1}{4}{\mathbb{E}}_{{\mathcal{D}}}\Big{[}R(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}+c\Big{\{}\beta\lambda^{2}+\varepsilon_{\operatorname{app}}({\mathcal{F}},\lambda)+\varepsilon_{\operatorname{gen}}({\mathcal{F}},n)\Big{\}},$
@@
${\mathbb{E}}_{{\mathcal{D}}}\Big{[}R(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}\leq{\mathbb{E}}_{{\mathcal{D}}}\Big{[}\widehat{R}_{{\mathcal{D}}}(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}+c^{\prime}\varepsilon_{\operatorname{gen}}({\mathcal{F}},n)$
$\leq\frac{1}{4}{\mathbb{E}}_{{\mathcal{D}}}\Big{[}R(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}+c\Big{\{}\beta\lambda^{2}+\varepsilon_{\operatorname{app}}({\mathcal{F}},\lambda)\Big{\}}$
$\quad+(c+c^{\prime})\varepsilon_{\operatorname{gen}}({\mathcal{F}},n).$
@@
$\|f-\phi\|_{L^{2}(\mu_{X})}\leq CN^{-s},$
$\|\nabla(f-\phi)\|_{L^{2}(\nu_{X})}\leq CN^{-(s-1)}.$
By using Lemma A.5 and set $\delta=1/n$, we find
$\log N(B_{0}n^{-1},{\mathcal{F}},L^{2}({\mathcal{D}}))\\ \lesssim LS\log(S)(\log n)\lesssim N^{d}\log^{2}N\log n.$ (54)
Substituting these estimates into Lemma V.2 yields
@@
$\lesssim\beta\lambda^{2}+CN^{-2s}+C\lambda N^{-2(s-1)}$
$\quad+C\log n\Big{(}\frac{N^{d}\log^{2}N\log n}{n}\Big{)}^{\frac{1}{2}}.$
Setting $N={\mathcal{O}}(n^{\frac{1}{d+4s}})$, and letting the regularization parameter be $\lambda={\mathcal{O}}(n^{-\frac{s}{d+4s}}\log^{2}n)$ deduce the desired result. ∎
D-B Semi-Supervised Deep Sobolev Regressor
Proof of Lemma V.5.
@@ D-B Semi-Supervised Deep Sobolev Regressor
$(u,v)_{L^{2}({\mathcal{S}})}=\frac{1}{m}\sum_{i=1}^{m}u(Z_{i})v(Z_{i}),~{}\|u\|_{L^{2}({\mathcal{S}})}^{2}=\frac{1}{m}\sum_{i=1}^{m}u^{2}(Z_{i}),$
for each $u,v\in L^{\infty}(\nu_{X})$. The proof is divided into four parts which are denoted by (I) to (IV):
@@
    By a same argument as (I) in the proof of Lemma V.2, we deduces
    ${\mathbb{E}}_{{\mathcal{S}}}\Big{[}\|\nabla(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0})\|_{L^{2}(\nu_{X})}^{2}-\|\nabla(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0})\|_{L^{2}({\mathcal{S}})}^{2}\Big{]}\\ \leq\varepsilon_{\operatorname{gen}}^{\operatorname{reg}}(\nabla{\mathcal{F}},m).$ (55)
    Here the generalization error $\varepsilon_{\operatorname{gen}}^{\operatorname{reg}}(\nabla{\mathcal{F}},m)$ associated to the regularization term are defined as
    $\varepsilon_{\operatorname{gen}}^{\operatorname{reg}}(\nabla{\mathcal{F}},m)\\ =B_{1}^{2}\inf_{\delta>0}\Big{\{}\max_{1\leq k\leq d}\frac{N(B_{1,k}\delta,D_{k}{\mathcal{F}},L^{2}({\mathcal{S}}))}{m}+\delta\Big{\}}.$
@@
    $-\lambda{\mathbb{E}}_{{\mathcal{S}}}\Big{[}(\nabla f_{0},\nabla(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f))_{L^{2}({\mathcal{S}})}\Big{]}\\ \leq\frac{1}{16}{\mathbb{E}}_{{\mathcal{S}}}\Big{[}R(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda})\Big{]}+c\Big{\{}\tilde{\beta}\lambda^{2}+\varepsilon_{\operatorname{gen}}^{\operatorname{reg}}(\nabla{\mathcal{F}},m)\Big{\}},$ (56)
@@
    $\lambda{\mathbb{E}}_{{\mathcal{D}},{\mathcal{S}}}\Big{[}\|\nabla(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0})\|_{L^{2}({\mathcal{S}})}^{2}\Big{]}+{\mathbb{E}}_{{\mathcal{D}},{\mathcal{S}}}\Big{[}\widehat{R}_{{\mathcal{D}}}(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda})\Big{]}$
@@
    $\quad+c\Big{\{}\tilde{\beta}\lambda^{2}+\varepsilon_{\operatorname{app}}({\mathcal{F}},\lambda)+\varepsilon_{\operatorname{gen}}^{\operatorname{reg}}(\nabla{\mathcal{F}},m)\Big{\}},$ (57)
@@
Step (I). By a same argument as Step (I) in the proof of Lemma V.2, we deduce the following inequality
${\mathbb{E}}_{{\mathcal{S}}}\Big{[}\|D_{k}(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0})\|_{L^{2}(\nu_{X})}^{2}-\|\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0}\|_{L^{2}({\mathcal{S}})}^{2}\Big{]}$
$\leq 4B_{1,k}^{2}\inf_{\delta>0}\Big{\{}\frac{N(B_{1,k}\delta,D_{k}{\mathcal{F}},L^{2}({\mathcal{S}}))}{m}+\delta\Big{\}},$
for each $1\leq k\leq d$. Summing over these equalities obtains (55) immediately.
@@
$={\mathbb{E}}_{{\mathcal{S}}}\Big{[}\sup_{f\in\operatorname{conv}({\mathcal{F}})}{\mathbb{E}}_{{\mathcal{S}}^{\prime}}\Big{[}\frac{1}{m}\sum_{i=1}^{m}D_{k}f_{0}(Z_{i}^{\prime})D_{k}f(Z_{i}^{\prime})\Big{]}$
$\quad-\frac{1}{m}\sum_{i=1}^{m}D_{k}f_{0}(Z_{i})D_{k}f(Z_{i})\Big{]}$
$\leq{\mathbb{E}}_{{\mathcal{S}}}{\mathbb{E}}_{{\mathcal{S}}^{\prime}}\Big{[}\sup_{f\in\operatorname{conv}({\mathcal{F}})}\frac{1}{m}\sum_{i=1}^{m}D_{k}f_{0}(Z_{i}^{\prime})D_{k}f(Z_{i}^{\prime})$
$\quad-\frac{1}{m}\sum_{i=1}^{m}D_{k}f_{0}(Z_{i})D_{k}f(Z_{i})\Big{]}$
$={\mathbb{E}}_{{\mathcal{S}}}{\mathbb{E}}_{{\mathcal{S}}^{\prime}}{\mathbb{E}}_{\varepsilon}\Big{[}\sup_{f\in\operatorname{conv}({\mathcal{F}})}\frac{1}{m}\sum_{i=1}^{m}\varepsilon_{i}\Big{(}D_{k}f_{0}(Z_{i}^{\prime})D_{k}f(Z_{i}^{\prime})$
$\quad-D_{k}f_{0}(Z_{i})D_{k}f(Z_{i})\Big{)}\Big{]}$
$={\mathbb{E}}_{{\mathcal{S}}}{\mathbb{E}}_{\varepsilon}\Big{[}\sup_{f\in\operatorname{conv}({\mathcal{F}})}\frac{1}{m}\sum_{i=1}^{m}\varepsilon_{i}D_{k}f_{0}(Z_{i})D_{k}f(Z_{i})\Big{]}$
$={\mathbb{E}}_{{\mathcal{S}}}{\mathbb{E}}_{\varepsilon}\Big{[}\sup_{f\in{\mathcal{F}}}\frac{1}{m}\sum_{i=1}^{m}\varepsilon_{i}D_{k}f_{0}(Z_{i})D_{k}f(Z_{i})\Big{]},$ (58)
where the second inequality follows from the Jensen’s inequality, and the last equality invokes the fact that the Rademacher complexity of the convex hull is equal to that of the original set. According to Ledoux-Talagrand contraction inequality [97, Lemma 5.7], we have
${\mathbb{E}}_{\varepsilon}\Big{[}\sup_{f\in{\mathcal{F}}}\frac{1}{m}\sum_{i=1}^{m}\varepsilon_{i}D_{k}f_{0}(Z_{i})D_{k}f(Z_{i})\Big{]}$ (59)
$\leq B_{1,k}{\mathbb{E}}_{\varepsilon}\Big{[}\sup_{f\in{\mathcal{F}}}\frac{1}{m}\sum_{i=1}^{m}\varepsilon_{i}D_{k}f(Z_{i})\Big{]}.$
Let $\delta>0$ and let $(D_{k}{\mathcal{F}})_{\delta}$ be an $L^{2}({\mathcal{S}})$ $(B_{1,k}\delta)$-cover of $D_{k}{\mathcal{F}}$. Suppose $|(D_{k}{\mathcal{F}})_{\delta}|=N(B_{1,k}\delta,D_{k}{\mathcal{F}},L^{2}({\mathcal{S}}))$. Then it follows from Cauchy-Schwarz inequality that for each $D_{k}f\in D_{k}{\mathcal{F}}$, there exists $(D_{k}f)_{\delta}\in(D_{k}{\mathcal{F}})_{\delta}$ such that
@@
$\leq\Big{(}\frac{1}{m}\sum_{i=1}^{m}\varepsilon_{i}^{2}\Big{)}^{1/2}\Big{(}\frac{1}{m}\sum_{i=1}^{m}(D_{k}f(Z_{i})-(D_{k}f)_{\delta}(Z_{i}))^{2}\Big{)}^{1/2}$
$\leq B_{1,k}\delta,$
@@
${\mathbb{E}}_{{\mathcal{S}}}{\mathbb{E}}_{\varepsilon}\Big{[}\sup_{D_{k}f\in{\mathcal{F}}}\frac{1}{m}\sum_{i=1}^{m}\varepsilon_{i}D_{k}f(Z_{i})\Big{]}$
$\leq{\mathbb{E}}_{{\mathcal{S}}}{\mathbb{E}}_{\varepsilon}\Big{[}\sup_{D_{k}f\in(D_{k}{\mathcal{F}})_{\delta}}\frac{1}{m}\sum_{i=1}^{m}\varepsilon_{i}D_{k}f(Z_{i})\Big{]}+B_{1,k}\delta$
$\leq B_{1,k}\Big{(}\frac{2\log N(B_{1,k}\delta,D_{k}{\mathcal{F}},L^{2}({\mathcal{S}}))}{m}\Big{)}^{1/2}+B_{1,k}\delta,$ (60)
where the last inequality holds from Massart’s lemma [97, Theroem 3.7]. Combining (58), (59) and (60) deduces